# Hyperbolic embedding of multilayer networks

### v2

@@ II.5 Layers extraction and radii attribution
where $r_{i}$ denotes the radial coordinate of node $i$, $\tanh$ is the hyperbolic tangent function, and $\beta$ is a scaling parameter. This formulation ensures that all nodes are embedded within the Poincaré disk, with nodes of degree zero mapped at the boundary ($r=1$) and the most central nodes positioned closer to the origin. The parameter $\beta$ controls the rate at which the radius decreases with increasing centrality and plays a role analogous to the popularity fading parameter in the original coalescent embedding method. A practical strategy for setting $\beta$ consists in prescribing the radius $r_{i}$ of a node $i$ with maximal (weighted) degree $w_{i}$, such that it lies within a desired distance from the origin (for instance, $r_{i}=0.1/0.2$) and solving $\beta=w_{i}/\text{arctanh}(r_{i})$. This ensures that node radii span a broad portion of the disk while maintaining an interpretable mapping between centrality and geometry.
@@ III.2 Nodes localization – An example on brain epilepsy networks
To illustrate the practical utility of our approach, we computed the hyperbolic Gaussian distribution of node positions (specifically from the left temporal lobes) within the Poincaré disk (See Sec. II.6). We evaluated two embedding strategies: our proposed multilayer method, in which each layer represents a different patient, and an independent embedding approach, wherein each patient is embedded separately and the resulting configurations are subsequently aligned via rotational transformation. The alignment is performed by rotating each embedding (beyond the reference layer) to minimize the sum of hyperbolic distances between corresponding nodes, effectively minimizing the $G_{\text{score}}$ defined Eq. 16. This procedure, similar to the one described in [37], ensures comparability across layers while preserving the geometric invariance of the embedding space.
@@
Figure 3: Hyperbolic embedding of patients with left temporal lobe epilepsy (red points) and controls (blue points). Visualization of the Gaussian distributions of node positions corresponding to the left temporal lobes (2 nodes per patients) within the Poincaré disk, using two embedding strategies: A) the proposed multilayer embedding approach, where each layer represents an individual patient; and B) an independent embedding of each patient followed by post hoc rotational alignment.

### v3

@@ II.5 Layers extraction and radii attribution
where $r_{i}$ denotes the radial coordinate of node $i$, $\tanh$ is the hyperbolic tangent function, and $\beta$ is a scaling parameter. This formulation ensures that all nodes are embedded within the Poincaré disk, with nodes of degree zero mapped at the boundary ($r=1$) and the most central nodes positioned closer to the origin. The parameter $\beta$ controls the rate at which the radius decreases with increasing centrality and plays a role analogous to the popularity fading parameter in the original coalescent embedding method.Unlike the original coalescent embedding, which assigns radii based on the degree ranking within a layer, here we directly use the (weighted) degree itself. This choice ensures that nodes with the same degree across different layers are mapped at consistent radial distances, while also naturally placing disconnected nodes ($w_{i}=0$) at the boundary ($r=1$). This adaptation makes the method more suitable for multilayer networks, where degree distributions can differ significantly between layers. A practical strategy for setting $\beta$ consists in prescribing the radius $r_{i}$ of a node $i$ with maximal (weighted) degree $w_{i}$, such that it lies within a desired distance from the origin (for instance, $r_{i}=0.1/0.2$) and solving $\beta=w_{i}/\text{arctanh}(r_{i})$. This ensures that node radii span a broad portion of the disk while maintaining an interpretable mapping between centrality and geometry.
@@ III.2 Nodes localization – An example on brain epilepsy networks
To illustrate the practical utility of our approach, we computed the hyperbolic Gaussian distribution of node positions (specifically from the left temporal lobes) within the Poincaré disk (See Sec. II.6). We evaluated two embedding strategies: our proposed multilayer method, in which each layer represents a different patient, and an independent embedding approach, wherein each patient is embedded separately and the resulting configurations are subsequently aligned via rotational transformation.For the rotational transformation, reflection symmetry (axis inversion) is also tested. The alignment error is computed both with and without reflection, and the configuration with the lowest error is retained.. The alignment is performed by rotating each embedding (beyond the reference layer) to minimize the sum of hyperbolic distances between corresponding nodes, effectively minimizing the $G_{\text{score}}$ defined Eq. 16. This procedure, similar to the one described in [37], ensures comparability across layers while preserving the geometric invariance of the embedding space.
@@
Figure 3: Hyperbolic embedding of patients with left temporal lobe epilepsy (red points) and controls (blue points). Visualization of the Gaussian distributions of node positions corresponding to the left temporal lobes (2 nodes per patients) within the Poincaré disk, using two embedding strategies: A) the proposed multilayer embedding approach, where each layer represents an individual patient $\beta=50,\mu=20$; and B) an independent embedding of each patient followed by post hoc rotational alignment.